OpenStudy (anonymous):

find the side labeled x and y round to two decimal places. i have to draw the triangle.

OpenStudy (anonymous):

do you have a picture to go along with this?

OpenStudy (anonymous):



OpenStudy (anonymous):

use SOHCAHTOA

OpenStudy (anonymous):

Do you know how to use that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright if you use sohcahtoa you can figure out the x and y that is needed for your answer! =)

OpenStudy (anonymous):

\[\cos 35 degree=\frac{ 2 }{ x }\]

OpenStudy (anonymous):

i just don't where to go from there?

OpenStudy (anonymous):

cos(35) = 2/x... all you gotta do is plug in cos35 into your calculator (if you can use it that is) then you will get a number... so it look like... # = 2/x then just solve for your x

OpenStudy (anonymous):

.

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

okay thanks

OpenStudy (unklerhaukus):

\[\cos(35°)=\frac 2x\] \[x=\frac 2{\cos(35°)}\]